OpenStudy (anonymous):

1/2 divided by 3= what? Medal to the best helper?

OpenStudy (solomonzelman):

\(\Large\color{black}{ \rm \frac{1}{2}\div3= \frac{1}{2\times 3}=? }\)

OpenStudy (anonymous):

Thank you

OpenStudy (solomonzelman):

Anytime !